Using Green's theorem evaluate the integral: ∫C (y³dx + 3x²dy) where C is the contour along the circle x² + y² = 1 taken counter clockwise.


1
Expert's answer
2021-06-09T17:54:23-0400

According to Green's theorem, "\int_C P dx + Q dy = \iint_D \left(\frac{ \partial Q}{\partial x} - \frac{\partial P}{\partial y}\right) dA".

For given vector field, "\frac{ \partial Q}{\partial x} - \frac{\partial P}{\partial y} = 6 x - 3 y^2", hence we need to evaluate surface integral"\iint (6 x - 3 y^2) dx dy" over the circle of radius 1, centered at the origin.

It is convenient to do it in polar coordinates, where surface element is "dS = r dr d\varphi", and "0\leq r \leq 1", "0\leq \varphi \leq 2 \pi":

"\int_0^{2\pi} d\varphi \int_0^1 r (6 r \cos \varphi - 3 r^2 \sin^2 \varphi) = \int_0^{2\pi}(\frac{6}{3}\cos \varphi - \frac{3}{4} \sin^2 \varphi)d \varphi ="

"\left[ 2 \sin \varphi - \frac{3}{4}(\frac{\varphi}{2} - \frac{1}{4} \sin 2 \varphi)\right]_0^{2 \pi} = -\frac{3}{4}\pi".